OpenStudy (anonymous):

factor the binomials completely. p^2 - 1/9

OpenStudy (anonymous):

same principle where (1/9) can be written as (1/3)^2 so = p^2 - (1/3)^2 = [p - (1/3)] * [p + (1/3)]

OpenStudy (anonymous):

ty so much